Project

General

Profile

Update 02 December 2016

Progress on Data Taking
Added by Kyberd, Paul 4 months ago

Dear All,

The Step IV data-taking programme with the lithium-hydride absorber continues to proceed according to plan. The present User Cycle, 2016/04, is dedicated to the collection of the data that will allow us to study the evolution of normalised transverse emittance in the Step IV lattice. Chris Rogers, Ao Liu and Jaroslaw Pasternak have carefully selected settings that should allow us to observe normalised-emittance reduction, cooling, over a range of input emittance and beam momentum.

To date two of our target of three settings have been recorded. Chis presented the attached plots at yesterday’s video conference. The plots show the distribution of transverse amplitude for samples of the data taken at 140 MeV/c with nominal input emittance of 6 mm and 10 mm. Good muons are defined to be those that pass through the whole apparatus, i.e. these muons are measured in both the upstream and in the downstream trackers. The green and blue histograms show the distribution of transverse amplitude measured in the upstream and the downstream tracker respectively. The red histograms show the distribution of transverse amplitude measured in the upstream tracker for those muons that are not observed after they have left the upstream tracking volume.

The red histogram indicates that transmission losses are dominated by muons with a large transverse amplitude upstream of the absorber. If the combination of energy loss in the LiH and the optics of the channel are such as to produce ionization cooling, then the population at low transverse amplitude measured in the downstream tracker should be larger than that measured in the upstream tracker. Looking closely at the plots, you can see an indication of such an effect in the (6,140) data. In the (10,140) data, an excess of events at low transverse amplitude is observed in the downstream distribution for transverse amplitudes below ~20 mm. As Chris says on his slide; MICE cools!

This is really an excellent start! We continue to take data in the 2016/04 1.2 configuration. On Monday we plan to reconfigure the magnetic channel to take data with a slightly smaller beta at the absorber. Analysis of the data will then allow us to study the cooling effect with lithium hydride.

If data taking continues with the efficiency that we have established over the past two Cycles, we shall be able to enjoy the Christmas break with lots of “data in the bank”!

With very best wishes …

Ken


Comments